# Explaining Large Language Models with gSMILE

### v4

@@ Abstract
Large Language Models (LLMs) such as GPT, LLaMA, and Claude achieve remarkable performance in text generation but remain opaque in their decision-making processes, limiting trust and accountability in high-stakes applications. We present gSMILE (generative SMILE), a model-agnostic, perturbation-based framework for token-level interpretability in LLMs. Extending the SMILE methodology, gSMILE uses controlled prompt perturbations, Wasserstein distance metrics, and weighted linear surrogates to identify input tokens with the most significant impact on the output. This process enables the generation of intuitive heatmaps that visually highlight influential tokens and reasoning paths. We evaluate gSMILE across leading LLMs (OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo, and Anthropic’s Claude 2.1) using attribution fidelity, attribution consistency, attribution stability, attribution faithfulness, and attribution accuracy as metrics. Results show that gSMILE delivers reliable human-aligned attributions, with Claude 2.1 excelling in attention fidelity and GPT-3.5 achieving the highest output consistency. These findings demonstrate gSMILE’s ability to balance model performance and interpretability, enabling more transparent and trustworthy AI systems.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 [1], Claude [2], and LLaMA [3] have rapidly become central to artificial intelligence, powering applications in healthcare, education, law, and scientific research [4, 5, 6, 7]. Built on the Transformer architecture [8] and trained on massive corpora, they generate fluent, context-aware responses for tasks ranging from dialogue and summarisation to reasoning and information extraction [6].
@@
Moreover, interpretability approaches must be model-agnostic. While Anthropic has recently showcased advanced interpretability research for Claude 3.5, using techniques such as circuit tracing and attribution graphs to visualise how the model plans, manipulates concepts, and sometimes even fabricates reasoning [12], such access to internal mechanisms is limited to others with privileged model access. For most practitioners and end users, leading LLMs like GPT and Claude are available only through APIs that restrict visibility into their parameters and architectures. As a result, practical interpretability techniques must be designed to explain model behaviour without relying on direct access to internal weights or structural details [6].
@@ 4 Proposed Method
Figure 4: SMILE flowchart for explaining image classification [13]
@@
Figure 5: gSMILE flowchart for Large Language Models
@@ 7 Conclusion
Interpretability remains a cornerstone of responsible AI, particularly in the context of large language models (LLMs) and instruction-based text generation. As AI systems become increasingly embedded in domains such as business, education, healthcare, and public services, it becomes essential for users to understand how these models generate outputs from their prompts. Without such insight, trust in AI systems diminishes, limiting adoption in real-world workflows. This underscores the need for transparent and interpretable solutions to demystify the behaviour of these black-box models.
Interpretability is not only a research priority but is likely to become a legal and operational requirement in alignment with emerging AI regulations and ethical standards. Organisations that fail to address interpretability may encounter regulatory hurdles and reduced user acceptance.
This work introduced a novel interpretability framework tailored for instruction-based text generation models. Leveraging gSMILE, we identified and visualised how individual words within a user prompt influence the model’s generated output. Our approach provides model-agnostic, local explanations by perturbing input prompts and analysing the resulting changes using Wasserstein-based distance metrics.
Whereas most existing interpretability methods highlight essential tokens in the model’s output [64, 65], gSMILE shifts attention to the input space. By attributing influence to specific words in the prompt, our method offers actionable insights into how linguistic phrasing shapes generation behaviour, a key consideration for prompt engineering and instruction design.
Our experiments show that gSMILE produces interpretable, stable, and faithful explanations across multiple instruction-tuned LLMs, including OpenAI GPT, Meta’s LLaMA 3.1 Instruct Turbo (70B), Claude.ai, and Gemma. The resulting heatmaps and attribution scores provide an intuitive understanding of model behaviour, empowering users to control and refine prompt design better. gSMILE was systematically evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including fidelity, stability, and consistency.
This work lays the groundwork for more transparent and trustworthy natural language systems. Future directions include extending gSMILE to support multi-turn dialogue, richer prompt hierarchies, and integration into AI safety, policy, and human-AI collaboration frameworks.
8 Future Works
@@
    Extension to Mathematical and Equation-Based Text Generation: Future work will explore the application of the proposed interpretability framework to models focused on symbolic reasoning and math-related tasks, such as equation solving, theorem generation, and formal derivation. Instruction-based mathematical models like Minerva and MathGLM require high precision in prompt design, and understanding how prompt components affect symbolic output will be essential in educational and scientific settings.
@@
    Application to Paraphrase Generation Models: The methodology can be extended to paraphrasing models, where semantic preservation and variation are central. Analysing how individual words influence paraphrased outputs’ style, tone, or fidelity can support better user control in writing assistance, content rephrasing, and translation tools.
@@
    Adaptation to Multi-Turn and Context-Aware Generation: Another direction is to adapt the framework for multi-turn and context-aware systems, including dialogue agents. This would involve tracking how prompt changes impact responses’ consistency, coherence, and relevance across conversational sequences.
  4. 4.
    Cross-Task Generalisation: While this work focuses on instruction-based text generation, the interpretability strategy can be generalised to broader NLP tasks, including summarisation, classification, and question answering. Evaluating how different model families respond to prompt perturbations can uncover model-specific behaviours and improve transparency.
  5. 5.
    Benchmarking Interpretability and Prompt Sensitivity: We plan to contribute to standardised evaluation protocols for prompt sensitivity, attribution robustness, and local fidelity. These benchmarks will support consistent assessment across models and help identify best practices for interpretable generation.
  6. 6.
    Optimisation of Computational Efficiency: To address the computational demands of gSMILE, future work will investigate dimensionality reduction techniques and approximation strategies. These approaches aim to reduce execution time while preserving the integrity and interpretability of the model’s outputs.
@@
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the gSMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/CELL_SMILE.

### v5

@@ Abstract
Large Language Models (LLMs) such as GPT, LLaMA, and Claude achieve remarkable performance in text generation but remain opaque in their decision-making processes, limiting trust and accountability in high-stakes applications. We present gSMILE (generative SMILE), a model-agnostic, perturbation-based framework for token-level interpretability in LLMs. Extending the SMILE methodology, gSMILE utilises controlled prompt perturbations, Wasserstein distance metrics, and weighted linear surrogates to identify input tokens that have the most significant impact on the output. This process enables the generation of intuitive heatmaps that visually highlight influential tokens and reasoning paths. We evaluate gSMILE across leading LLMs (OpenAI’s gpt-3.5-turbo-instruct, LLaMA 3.1 Instruct Turbo, and Anthropic’s Claude 2.1) using attribution metrics such as fidelity, consistency, stability, faithfulness, and accuracy. Results show that gSMILE delivers reliable human-aligned attributions, with Claude 2.1 excelling in attention fidelity and GPT-3.5 achieving the highest output consistency. These findings demonstrate gSMILE’s ability to strike a balance between model performance and interpretability, enabling more transparent and trustworthy AI systems.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT‑4 [1], Claude [2], and LLaMA [3] have rapidly become central to artificial intelligence, powering applications in healthcare, education, law, and scientific research [4, 5, 6, 7]. Built on the Transformer architecture [8] and trained on massive corpora, they generate fluent, context‑aware responses for tasks ranging from dialogue and summarisation to reasoning and information extraction [6].
@@
Moreover, interpretability approaches must be model‑agnostic. While Anthropic has recently showcased advanced interpretability research for Claude 3.5, using techniques such as circuit tracing and attribution graphs to visualise how the model plans, manipulates concepts, and sometimes even fabricates reasoning [12], such access to internal mechanisms is limited to others with privileged model access. For most practitioners and end users, leading LLMs like GPT and Claude are available only through APIs that restrict visibility into their parameters and architectures. As a result, practical interpretability techniques must be designed to explain model behaviour without relying on direct access to internal weights or structural details [6].
@@ 4 Proposed Method
Figure 4: SMILE flowchart for explaining image classification [13]
@@ 4.1 Text Generating and Perturbed Prompts
Figure 5: gSMILE flowchart for Large Language Models
@@ 5.1.1 Input Prompt Explainability
5.1.2 Gender Fairness via Token-Level Attributions
We further examined whether gendered phrasing affects attribution patterns by applying gSMILE to three closely related prompts: “What is the meaning of life?”, “What is the meaning of life from men’s perspective?”, and “What is the meaning of life from women’s perspective?”. In all three models, the neutral formulation consistently highlighted the tokens meaning and life, underscoring their central role in shaping the response. Once a gender qualifier was introduced, however, the attribution heatmaps shifted, revealing model-specific sensitivities. As shown in Fig. 9, GPT-3.5 tended to assign greater weight to the token women than men, suggesting an asymmetric influence of gender terms. Claude 3.5 distributed its attention more evenly, balancing the contribution of both gendered tokens while still raising their importance relative to the baseline. By contrast, LLaMA 3.1 exhibited higher variability, with attribution sometimes amplifying one gender reference more strongly than the other.
Figure 9: Comparison of gSMILE token-level attributions for GPT-3.5, LLaMA 3.1, and Claude 3.5 under three prompts: neutral, “men perspective”, and “women perspective”. Colour intensity encodes the relative influence of each token on the output distribution.
These findings illustrate how gSMILE exposes fairness-related sensitivities that remain hidden if only the generated text is inspected. A gender term that absorbs disproportionate attribution suggests an underlying association that could bias the model’s reasoning. From a practical standpoint, this analysis highlights that avoiding gendered modifiers in prompts helps maintain attention on the semantic core of the task, thereby reducing the risk of unintended attribution drift.
@@ 6 Limitations
Taken together, these limitations highlight the trade-off between generality and practicality in perturbation-based interpretability. They also point toward future work in designing adaptive, cost-efficient perturbation strategies and incorporating causal verification mechanisms to strengthen the robustness of explanations.
@@ 7 Conclusion
This work introduced gSMILE, a generative, model-agnostic framework for explaining LLM behaviour through token-level attribution and visual reasoning aids. Whereas most existing interpretability methods highlight essential tokens in the model’s output [64, 65], gSMILE shifts attention to the input space. By integrating perturbation analysis with Wasserstein-based distance measures and weighted linear surrogates, gSMILE offers a principled approach to local interpretability without requiring access to internal model parameters.
Our evaluation across three instruction-tuned LLMs revealed clear performance contrasts:
  • Claude 2.1 achieved the highest attribution fidelity (AttAUROC = 0.88) and lowest error measures, indicating more semantically focused attention.
  • GPT-3.5-turbo-instruct exhibited exceptional consistency (lowest variance and standard deviation across repeated runs), making it highly predictable in explanation output.
  • LLaMA 3.1 performed competitively but showed greater variability in token attribution, suggesting sensitivity to prompt phrasing.
Overall, the resulting heatmaps and attribution scores provide an intuitive understanding of model behaviour, empowering users to control and refine prompt design better.
gSMILE was evaluated and compared against existing methods (LIME, Bay-LIME) using multiple explainability metrics, including attribution metrics such as fidelity, consistency, stability, faithfulness, and accuracy. We systematically assessed how well each method captured token-level influence in LLMs. The results show that gSMILE consistently achieved higher fidelity in approximating black-box behaviour, while also producing more stable and repeatable attributions under small prompt perturbations. Compared to LIME and Bay-LIME, gSMILE demonstrated improved robustness and alignment with human-interpretable patterns, underscoring its suitability for practical use in settings where reliable interpretability is critical.
These findings justify gSMILE’s contribution in two ways. First, it demonstrates that explainability metrics can meaningfully differentiate between LLMs beyond standard benchmark accuracy, providing new insight into their operational reliability. Second, it shows that a purely black-box, post-hoc approach can produce explanations with stability and fidelity comparable to privileged-access interpretability tools.
At the same time, the limitations identified (prompt sensitivity, hallucination effects, computational overhead, and deployment constraints) highlight that the path toward fully practical interpretability remains open. Addressing these will be central to future research. Looking ahead, we see three promising directions:
@@
    Adaptive perturbation strategies: Future work can move beyond random sampling by using targeted perturbations that focus only on influential tokens or phrases. This reduces computational overhead while preserving fidelity. Inspired by approaches such as IBM’s CELL framework, perturbations could involve changing specific words rather than only removing or keeping them. This makes it possible to highlight the importance of selected words in longer sentences efficiently.
@@
    Causal verification methods: To mitigate hallucination-driven or spurious attributions, causal analysis can be integrated (e.g., perturbing tokens while holding other context fixed) to confirm whether an attribution genuinely drives the model’s behaviour. This would strengthen robustness in high-stakes applications.
@@
    Extension to multimodal and multilingual models: gSMILE currently focuses on English text. Extending the framework to multilingual LLMs and multimodal systems (text–image or text–speech) would test whether stability and fidelity hold across diverse input spaces, improving generalisability for real-world deployment.
By uniting rigorous evaluation with interpretable visualisations, gSMILE makes LLMs more transparent without privileged access to their internals, thereby providing both researchers and practitioners with a practical tool for building more trustworthy AI systems.
@@
The datasets used in this article are publicly available, and the code supporting this paper is published online on GitHub. For the gSMILE explainability framework, please refer to Dependable-Intelligent-Systems-Lab/xwhy. For the proposed algorithm introduced in this paper, see Sara068/LLM_SMILE.